# Unusual low-temperature behavior in the half-filled band of the one-dimensional extended Hubbard model in atomic limit

### v1

@@ Acknowledgements.
O. R. and S. M. S. thank CNPq and FAPEMIG for partial financial support. M. L. L. and M. S. S. Pereira also acknowledge the financial support of FAPEAL. O. D. acknowledges support through the EURIZON project (project #3025), which is funded by the European Union under grant agreement No.871072.
@@ Appendix A Transfer matrix around pseudo-transition
The phenomenon of pseudo-transition arises near the alternating pairs $AP$ and paramagnetic $PM$ phases. Although the solution of (7) is exact and analytic, the analytical expression for finding the pseudo-critical temperature could becomes a cumbersome task due to involving cubic root expression. An alternative approach to determine this condition is to split the transfer matrix into two terms, namely, $\mathbf{V}=\mathbf{V}_{0}+\varsigma\mathbf{V}_{1}$. Here, $\mathbf{V}_{0}$ represents the core structure matrix, whose matrix elements just includes the half-filled case, while $\varsigma\mathbf{V}_{1}$ describes a small perturbation applied to the matrix, corresponding to non-half-filled band. The detailed formulations of these matrices are provided as follows:
@@
From this solution one can clearly verify that $v_{1}^{(0)}>v_{2}^{(0)}$, however, we cannot affirm any condition between $v_{0}^{(0)}$ and $v_{1}^{(0)}$, since there is no restriction, so one can establish the following condition $v_{1}^{(0)}=v_{0}^{(0)}$ this would be essential to find the pseudo-critical temperature, as discussed in Sec. III.3. The corresponding eigenvectors can be expressed by
@@
It is evident that the eigenvalues up to second-order correction indicate that $\lambda_{1}>\lambda_{2}$. However, we cannot affirm just by looking the perturbative eigenvalues which eigenvalues is the largest one $\lambda_{0}$ or $\lambda_{1}$. Although we have confirmed at the end of Sec. II.1 that $\lambda_{0}$ must be the largest one. Nevertheless, our perturbative result could be useful to find the condition of pseudo-critical temperature up to second-order approximation. Therefore, by using the condition given in (33) we can find the following relation
@@
this result would be relevant when we look for a more accurate value of the pseudo-critical temperature.

### v2

@@ Acknowledgements.
O. R. and S. M. S. thank CNPq and FAPEMIG for partial financial support. M. L. L. and M. S. S. Pereira also acknowledge the financial support of FAPEAL. O. D. acknowledges support through the EURIZON project (project #3025), which is funded by the European Union under grant agreement No.871072. L.M.V acknowledges the support of Donostia International Physics Center and iKur estrategia.
@@ Appendix A Transfer matrix around pseudo-transition
The phenomenon of pseudo-transition arises near the alternating pairs $AP$ and paramagnetic $PM$ phases. Although the solution of (7) is exact and analytic, the analytical expression for finding the peculiar temperature could becomes a cumbersome task due to involving cubic root expression. An alternative approach to determine this condition is to split the transfer matrix into two terms, namely, $\mathbf{V}=\mathbf{V}_{0}+\varsigma\mathbf{V}_{1}$. Here, $\mathbf{V}_{0}$ represents the core structure matrix, whose matrix elements just includes the half-filled case, while $\varsigma\mathbf{V}_{1}$ describes a small perturbation applied to the matrix, corresponding to non-half-filled band. The detailed formulations of these matrices are provided as follows:
@@
From this solution one can clearly verify that $v_{1}^{(0)}>v_{2}^{(0)}$, however, we cannot affirm any condition between $v_{0}^{(0)}$ and $v_{1}^{(0)}$, since there is no restriction, so one can establish the following condition $v_{1}^{(0)}=v_{0}^{(0)}$ this would be essential to find the peculiar temperature, as discussed in Sec. III.3. The corresponding eigenvectors can be expressed by
@@
It is evident that the eigenvalues up to second-order correction indicate that $\lambda_{1}>\lambda_{2}$. However, we cannot affirm just by looking the perturbative eigenvalues which eigenvalues is the largest one $\lambda_{0}$ or $\lambda_{1}$. Although we have confirmed at the end of Sec. II.1 that $\lambda_{0}$ must be the largest one. Nevertheless, our perturbative result could be useful to find the condition of peculiar temperature up to second-order approximation. Therefore, by using the condition given in (33) we can find the following relation
@@
this result would be relevant when we look for a more accurate value of the peculiar temperature.